OpenStudy (anonymous):

My question has two parts (a) in standard form and (b) in slope-intercept form for the line described.. through (3,-2), parallell to 2x-y=5

OpenStudy (anonymous):

so the line it's parallel to is y=2x-5, so take the slope (2) In point slope form, you then have y+2=2(x-3), which is a) 2x-y=8 b) y=2x-8